# DRAT Proofs of Unsatisfiability for SAT Modulo Monotonic Theories

### 19 Jan 2024 v1

@@ 0.A.4 Proof of Lemma 3 (Instantiation-Based Horn Upper-Bound)
For the forward direction, if $\mathcal{M}_{A}\Rightarrow\bar{\mathbf{p}}$ is a theory lemma, then for every complete assignment $M$ over variables $A$ that are extended from $\mathcal{M}_{A}$, it must be the case that $p(M(A))=\bot$. Therefore, $\Sigma_{\mathbf{p}}^{+}|_{\bar{\mathbf{p}}\cup M}$ is satisfiable since $\Sigma_{\mathbf{p}}^{+}$ is a propositional definition of $\mathbf{p}$. Let $\mathcal{M}_{A}^{+}$ be the complete assignment on $A$ by assigning $\top$ to every unassigned variable in $A$, and let $\mathcal{M}_{X}$ be the partial solution to the formula $\Sigma_{\mathbf{p}}^{+}|_{\bar{\mathbf{p}}\cup\mathcal{M}_{A}^{+}}$ on the auxiliary variables $X$. Since $\Sigma_{\mathbf{p}}^{+}$ is a monotonic definition which contains only the negative atoms from $A$, the solution $\mathcal{M}_{X}$ to $\Sigma_{\mathbf{p}}^{+}|_{\bar{\mathbf{p}}\cup\mathcal{M}_{A}^{+}}$ is also a solution to $\Sigma_{\mathbf{p}}^{+}|_{\bar{\mathbf{p}}\cup M}$ for every complete assignment $M$ over variables $A$ that are extended from $\mathcal{M}_{A}$. Therefore, the formula $\Sigma_{\mathbf{p}}^{+}|_{\bar{\mathbf{p}}\cup\mathcal{M}_{X}}\Rightarrow\bar{\mathbf{p}}$ logically implies $M\Rightarrow\bar{\mathbf{p}}$ for every $M$ extending from $\mathcal{M}_{A}$. Therefore, $\mathcal{M}_{A}\Rightarrow\bar{\mathbf{p}}$ is also a logical consequence of $(\Sigma_{\mathbf{p}}^{+}|_{\bar{\mathbf{p}}\cup\mathcal{M}_{X}})\Rightarrow\bar{\mathbf{p}}$. By Lemma. 3, the formula $(\Sigma{}_{\mathbf{p}}^{+}|_{\bar{\mathbf{p}}\cup\mathcal{M}_{X}}\Rightarrow\bar{\mathbf{p}})$ is Horn, and thus $(\Sigma_{\mathbf{p}}^{+}|_{\bar{\mathbf{p}}\cup\mathcal{M}_{X}}\Rightarrow\bar{\mathbf{p}})\vdash_{1}(\mathcal{M}_{A}\Rightarrow\bar{\mathbf{p}})$. ∎
@@
For the backward direction, suppose there exists some witness $\mathcal{M}_{X}$ such that $((\Sigma_{\mathbf{p}}^{+}|_{\bar{\mathbf{p}}\cup\mathcal{M}_{X}})\Rightarrow\bar{\mathbf{p}})\vdash_{1}(\mathcal{M}_{A}\Rightarrow\bar{\mathbf{p}})$. By Lemma 3, $\Sigma_{\mathbf{p}}^{+}|_{\bar{\mathbf{p}}\cup\mathcal{M}_{X}}$ is a Horn upper-bound of any propositional definition of $\bar{\mathbf{p}}$. Therefore, $\mathcal{M}_{A}\Rightarrow\bar{\mathbf{p}}$ is a valid theory lemma. ∎
Appendix 0.B Workflow of Horn Approximation for Verifying Theory Lemmas
The workflow of constructing proof-specific Horn definition for theory predicates is shown in Fig. 4.
@@ 0.C.3 Maximum Flow
Appendix 0.D Appendix: Full Experimental Results Table

### 04 Apr 2024 v2

@@ 0.A.4 Proof of Lemma 3 (Instantiation-Based Horn Upper-Bound)
For the forward direction, if $\mathcal{M}_{A}\Rightarrow\overline{\mathbf{p}}$ is in the theory of $\overline{\mathbf{p}}$, then for every complete assignment $M$ over variables $A$ that are extended from $\mathcal{M}_{A}$, it must be the case that $p(M(A))=\bot$. Therefore, since $\Sigma_{\mathbf{p}}^{+}$ is a propositional definition of $\mathbf{p}$ (Def. 2), we know that $\Sigma_{\mathbf{p}}^{+}\not\models(M\Rightarrow\mathbf{p})$. This means that $\Sigma_{\mathbf{p}}^{+}\wedge\neg(M\Rightarrow\mathbf{p})$ is satisfiable; equivalently, that $\Sigma_{\mathbf{p}}^{+}\wedge M\wedge\overline{\mathbf{p}}$ is satisfiable. Therefore, $\Sigma_{\mathbf{p}}^{+}|_{\overline{\mathbf{p}}\cup M}$ must also be satisfiable.
Now, let us consider a specific complete assignment to the variables in $A$, based on the given $\mathcal{M}_{A}$: let $\mathcal{M}_{A}^{+}$ be the complete assignment on $A$ by assigning $\top$ to every unassigned variable in $A$, and let $\mathcal{M}_{X}$ be a satisfying assignment (on the auxiliary variables $X$) to the formula $\Sigma_{\mathbf{p}}^{+}|_{\overline{\mathbf{p}}\cup\mathcal{M}_{A}^{+}}$, which we know is satisfiable from the preceding paragraph.
Since $\Sigma_{\mathbf{p}}^{+}$ is a monotonic definition (Def. 5), it can contain only negative atoms from $A$, so the solution $\mathcal{M}_{X}$ to $\Sigma_{\mathbf{p}}^{+}|_{\overline{\mathbf{p}}\cup\mathcal{M}_{A}^{+}}$ is also a solution to $\Sigma_{\mathbf{p}}^{+}|_{\overline{\mathbf{p}}\cup M}$ for every complete assignment $M$ over variables $A$ that is extended from $\mathcal{M}_{A}$. Thus, we have established the claim (1) that $\Sigma_{\mathbf{p}}^{+}|_{\overline{\mathbf{p}}\cup\mathcal{M}_{X}\cup\mathcal{M}_{A}}$ is satisfiable.
Moving on to claim (2) that $(\Sigma_{\mathbf{p}}^{+}|_{\overline{\mathbf{p}}\cup\mathcal{M}_{X}}\Rightarrow\overline{\mathbf{p}})\vdash_{1}(\mathcal{M}_{A}\Rightarrow\overline{\mathbf{p}})$, the key fact that we have just established is that $\mathcal{M}_{X}$ is a solution to $\Sigma_{\mathbf{p}}^{+}|_{\overline{\mathbf{p}}\cup M}$ for every complete assignment $M$ over variables $A$ that is extended from $\mathcal{M}_{A}$; in other words, we know that $\Sigma_{\mathbf{p}}^{+}|_{\overline{\mathbf{p}}\cup M\cup\mathcal{M}_{X}}=\top$ for all complete assignments $M$ over variables $A$ that are extended from $\mathcal{M}_{A}$. We will return to this fact shortly.
Now, we will prove that $(\Sigma_{\mathbf{p}}^{+}|_{\overline{\mathbf{p}}\cup\mathcal{M}_{X}}\Rightarrow\overline{\mathbf{p}})\models(\mathcal{M}_{A}\Rightarrow\overline{\mathbf{p}})$ by contradiction. Suppose
$(\Sigma_{\mathbf{p}}^{+}|_{\overline{\mathbf{p}}\cup\mathcal{M}_{X}}\Rightarrow\overline{\mathbf{p}})\not\models(\mathcal{M}_{A}\Rightarrow\overline{\mathbf{p}}).$
Manipulating the expression:
$(\Sigma_{\mathbf{p}}^{+}|_{\overline{\mathbf{p}}\cup\mathcal{M}_{X}}\Rightarrow\overline{\mathbf{p}})\not\models(\mathcal{M}_{A}\Rightarrow\overline{\mathbf{p}})$
iff $(\Sigma_{\mathbf{p}}^{+}|_{\overline{\mathbf{p}}\cup\mathcal{M}_{X}}\Rightarrow\overline{\mathbf{p}})\wedge\neg(\mathcal{M}_{A}\Rightarrow\overline{\mathbf{p}})\mbox{\ is satisfiable.}$.
iff $(\Sigma_{\mathbf{p}}^{+}|_{\overline{\mathbf{p}}\cup\mathcal{M}_{X}}\Rightarrow\overline{\mathbf{p}})\wedge\mathcal{M}_{A}\wedge\mathbf{p}\mbox{\ is satisfiable.}$.
iff ($\left(\neg(\Sigma_{\mathbf{p}}^{+}|_{\overline{\mathbf{p}}\cup\mathcal{M}_{X}})\vee\overline{\mathbf{p}}\right)\wedge\mathcal{M}_{A}\wedge\mathbf{p}\mbox{\ is satisfiable.}$.
iff $\neg(\Sigma_{\mathbf{p}}^{+}|_{\overline{\mathbf{p}}\cup\mathcal{M}_{X}})\wedge\mathcal{M}_{A}\wedge\mathbf{p}\mbox{\ is satisfiable.}$.
This means there must exist a complete assignment $M^{\prime}$ over variables $A$ that is extended from $\mathcal{M}_{A}$, such that $\neg(\Sigma_{\mathbf{p}}^{+}|_{\overline{\mathbf{p}}\cup\mathcal{M}_{X}})|_{M^{\prime}}=\top$, or equivalently, that $(\Sigma_{\mathbf{p}}^{+}|_{\overline{\mathbf{p}}\cup\mathcal{M}_{X}\cup M^{\prime}})=\bot$. But this contradicts the fact that $\Sigma_{\mathbf{p}}^{+}|_{\overline{\mathbf{p}}\cup M\cup\mathcal{M}_{X}}=\top$ for all complete assignments $M$ over variables $A$ that are extended from $\mathcal{M}_{A}$. Therefore, $(\Sigma_{\mathbf{p}}^{+}|_{\overline{\mathbf{p}}\cup\mathcal{M}_{X}}\Rightarrow\overline{\mathbf{p}})\models(\mathcal{M}_{A}\Rightarrow\overline{\mathbf{p}})$.
By Lemma 3, we construct $(\Sigma_{\mathbf{p}}^{+}|_{\overline{\mathbf{p}}\cup\mathcal{M}_{X}}\Rightarrow\overline{\mathbf{p}})$ as (dual) Horn clauses, and thus $(\Sigma_{\mathbf{p}}^{+}|_{\overline{\mathbf{p}}\cup\mathcal{M}_{X}}\Rightarrow\overline{\mathbf{p}})\vdash_{1}(\mathcal{M}_{A}\Rightarrow\overline{\mathbf{p}})$. This completes the proof in the forward direction.
@@
For the backward direction, suppose there exists some witness $\mathcal{M}_{X}$ such that $(\Sigma_{\mathbf{p}}^{+}|_{\overline{\mathbf{p}}\cup\mathcal{M}_{X}}\Rightarrow\overline{\mathbf{p}})\vdash_{1}(\mathcal{M}_{A}\Rightarrow\overline{\mathbf{p}})$. By Lemma 3, $(\Sigma_{\mathbf{p}}^{+}|_{\overline{\mathbf{p}}\cup\mathcal{M}_{X}}\Rightarrow\overline{\mathbf{p}})$ is a Horn upper-bound of any propositional definition $\Sigma_{\overline{\mathbf{p}}}$ of $\overline{\mathbf{p}}$. Therefore,
$\Sigma_{\overline{\mathbf{p}}}$ $\models$ $(\Sigma_{\mathbf{p}}^{+}|_{\overline{\mathbf{p}}\cup\mathcal{M}_{X}}\Rightarrow\overline{\mathbf{p}})$
$\vdash_{1}$ $\mathcal{M}_{A}\Rightarrow\overline{\mathbf{p}},$
so $\mathcal{M}_{A}\Rightarrow\overline{\mathbf{p}}$ is a valid theory lemma. ∎
Appendix 0.B Workflow for Verifying Theory Lemmas
@@ 0.C.3 Maximum Flow
Appendix 0.D Full Experimental Results Table